OpenStudy (anonymous):

Solve for the variable: -12x + 12 = -24 A. -2 B. 2 C. 3 D. 4

OpenStudy (jonnyvonny):

When solving for y, or any variable, we must do reverse PEMDAS. However, to make this A LOT simpler, we may divide both sides of the equation by a certain number; what is it?

OpenStudy (anonymous):

-12x+12=-24 -12x=-36 x=3 C

OpenStudy (jonnyvonny):

Good job, you tool.

OpenStudy (anonymous):

thanks guy

OpenStudy (anonymous):

Thanks!